OpenStudy (anonymous):

Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution.

OpenStudy (anonymous):

Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution. Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions. Part 3: Provide a system of TWO equations in slope-intercept form with infinitely many solutions. Using complete sentences, explain why this system has infinitely many solutions.

OpenStudy (anonymous):

y = x +7 y = 2x +3 There is only one solution because (4,11) is the only point where two lines intersect.

OpenStudy (anonymous):

part 2 ?

OpenStudy (anonymous):

y = x + 7 y = x + 3 There is no solution because there is no point where the two lines intersect.

OpenStudy (anonymous):

And for part three, and equations that are the same have infinite solutions.

OpenStudy (anonymous):

how would you write that out because i need to explain each one

OpenStudy (radar):

While y= x + 7 and y = 2x + 3 has only one solution, i believe the intersection is at (4,3) rather than (4, 11) Wouldn't you agree?

OpenStudy (radar):

The part 3, could use these: y = x + 7 2y=2x + 14 They are identities and have infinite solutions.

OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9 HELP ME PLEASE.

OpenStudy (radar):

Part 1 The solution is not correct. x = 2y + 14 x = -3y + 9 we know that x = x so we can write: 2y + 14 = -3y + 9 5y = -5 y = -1 substituting the -1 value into one of the original equation solving for x. x = -2 +14 x=12